solve the following recurrence relations.


a) T(n)=T(n-1)+1 , for n≥ 2 and T(1)=1.


b) "a_{n+1}-2a_n=2n," for n≥1 and T(1)=2.


1
Expert's answer
2022-05-11T14:35:51-0400

a). We see that "T(2)=T(1)+1=2;" "T(3)=T(2)+1=2+1=3". We make a conjecture that "T(k)=k", "k\in{\mathbb{N}}". Then, "T(k+1)=T(k)+1=k+1". Thus, the conjecture is true.

b). "a_{n+1}=2(a_n+n)=2(2(a_{n-1}+n-1)+n)=2^2a_{n-1}+2^2(n-1)+2n=2^3a_{n-2}+2^3(n-2)+2^2(n-1)+2n"

In case we continue the procedure, we will get: "a_{n+1}=2^{n}2+2^{n}+2^{n-1}2+2^{n-2}3+..+2n". We can present part of a formula as: "2^{n}+2^{n-1}2+2^{n-2}3+..+2n=2^n(1+2\cdot2^{-1}+3\cdot2^{-2}+\ldots+n2^{1-n})=2^n\sum_{k=1}^nk2^{1-k}=2^n(-4(\frac12)^{n+1}(n+1)-4(\frac12)^{n+1}+4)=-2(n+2)+2^{n+2}"

We used a formula for arithmetic-geometric progression.

We receive: "a_n=2^n-2(n+1)+2^{n+1}". It remains to check the obtained formula for "a_n" : "2a_n+2n=2^{n+1}-4(n+1)+2^{n+2}+2n=2^{n+1}-2(n+2)+2^{n+2}". Thus, it is true.

Answer: a). "T(k)=k", "k\in{\mathbb{N}}". b). "a_{n}=2^n-2(n+1)+2^{n-1}", "n\in{\mathbb{N}}.".